A point charge of magnitude q is at the center of a cube with sides of length L

Part A)

What is the electric flux ? through each of the six faces of the cube?

Use ?0 for the permittivity of free space.

Part B)

What would be the flux ?1 through a face of the cube if its sides were of length L1 ?

Use ?0 for the permittivity of free space.
Answer #1

Gauss law states thatthe electric flux through a closed surface is equal to ^ times the charge enclosed by the surface. Thus,
Answer #2

Electric flux is given by:

\phi=\frac{q}{\varepsilon _o}

Part A

\Rightarrow \phi=\frac{q}{\varepsilon _o}

this is the total flux through the cube

since a cube has six sides therefore flux through each face is given by:

\Rightarrow \phi_s=\frac{\phi}{6}=\frac{q}{6\varepsilon _o}

\Rightarrow \phi_s=\frac{q}{6\varepsilon _o}

Part B

if its sides were of length L1 then also the magnitude of the flux will remain same as total charge enclosed is same

therefore now flux is given by

\Rightarrow \phi_s=\frac{q}{6\varepsilon _o}